OpenStudy (anonymous):

Suppose y varies directly with x and y = 72 when x = 6. What direct variation equation relates x and y? What is the value of y when x = 10?

OpenStudy (freckles):

y varies directly with x mean for some constant k with have the equation y=kx

OpenStudy (freckles):

you are given a point on the line y=kx that point being (6,72) <---plug into find the constant k

OpenStudy (anonymous):

thank you

OpenStudy (freckles):

let me know when you have found the constant k

OpenStudy (anonymous):

no sorry

OpenStudy (freckles):

so you aren't able to solve the equation 72=6k for k?

OpenStudy (freckles):

if you aren't able to, can you at least think of number when multiplied to 6 gives you 72?

OpenStudy (anonymous):

A. y=1/9x;10/9 B. y=x+9;19 C. y=9x;90 D. y=x−9;1

OpenStudy (freckles):

none of those are the answer

OpenStudy (freckles):

but anyways back to the question

OpenStudy (freckles):

can you solve 72=6k for k?

OpenStudy (anonymous):

i dont really know

OpenStudy (freckles):

you must have solved linear equations like 6k=72 before? notice the k is being multiplied by 6 divide both sides by 6 to isolate k

OpenStudy (anonymous):

k=12

OpenStudy (freckles):

great so this is what we have done so far we were given that y varies directly with x which means the equation has form y=kx where k is a constant we we given that y=72 when x=6 which means the point (6,72) is on the line y=kx we plug this point in giving us 72=k*6 or 72=6k and then divided both sides by 6 72/6=k 12=k so we now know the constant k for any point (x,y) for this line which is 12 so the equation is y=12x now we are asked to find y when x is 10 so replace x with 10 now to find the corresponding y value

OpenStudy (anonymous):

okay

OpenStudy (freckles):

did you want to tell me what you got for y when you replaced x with 10?